Calculate the entropy change for 3 kg of ice melting. Given Hf = 333.5 kJ/kgH20


1
Expert's answer
2022-01-21T14:40:08-0500

"S=\frac{mH}T=\frac{3\cdot 333500}{273}=3.7~\frac{kJ}K."